an electron has a speed of 500m/s accurate to 0.0001% with what fundamental accuracy can we locate the position of the electron?

1
Expert's answer
2022-03-26T15:31:30-0400

"\Delta x\Delta p=\frac h2,"

"\Delta x=\frac h{2m\Delta v}=7.2\cdot 10^{-7}~m."